OpenStudy (anonymous):

find a simple formula for the nth term for the following: 1,3,7,15,31,63,...

OpenStudy (helder_edwin):

\[ \large 1=2^1-1 \] \[ \large 3=2^2-1 \] \[ \large 7=2^3-1 \]

OpenStudy (anonymous):

so that would make \[a _{n}= 2^{n}-1\]

OpenStudy (helder_edwin):

yes